Perpendicular to y= 1/3x + 2; (3,3)
Answer by mananth(16946)

1 y = 1/3 x + 2
Divide by 1
y = 1/ 3 x + 2
Compare this equation with y=mx+b
slope m = 1/ 3

The slope of a line perpendicular to the above line will be the negative reciprocal -3
m1*m2=-1


m= -3 ,point ( 3 , 3 )
Find b by plugging the values of m & the point in
y=mx+b
3 = -9 + b
b= 12
m= -3
The required equation isy =-3x + 12
m.ananth@hotmail.ca